OpenStudy (tocuteforyou2):

The square root of a number is between 8 and 9. Which of these could be that number? Question 3 options: A) 17 B) 63 C) 71 D) 89

zepdrix (zepdrix):

8*8 = 64. So we know that \(\rm \sqrt{64}\) is \(\rm 8\). This another way we can write the number 8. Similarly, \(\rm 9\) can be written as \(\rm \sqrt{81}\). So when you're looking for a number between 8 and 9, \(\large\rm 8<x<9\) You're really looking for a number between \(\rm \sqrt{64}\) and \(\rm \sqrt{81}\). \(\large\rm \sqrt{64}<x<\sqrt{81}\)

zepdrix (zepdrix):

So for example: \(\rm \sqrt{64}<\sqrt{80}<\sqrt{81}\) Square root of the number 80 would lie between 8 and 9.

zepdrix (zepdrix):

Think you can figure it out from there? :d

OpenStudy (tocuteforyou2):

no thats why i asked for help lol which is it A B C D

zepdrix (zepdrix):

I'm not going to give you a direct answer :) I'm trying to help you learn it. That's the point of all that typing :P

OpenStudy (tocuteforyou2):

oh ok its 71?

zepdrix (zepdrix):

71 is between 64 and 81, so yayyyy that one works out \c:/

zepdrix (zepdrix):

\(\large\rm \color{indianred}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (tocuteforyou2):

thanks

OpenStudy (tocuteforyou2):

i have another question

OpenStudy (thefatdog):

What is it?